OpenStudy (anonymous):

what is the slope between the two points (5, 1) and (4, 6)?

OpenStudy (anonymous):

Well, we have to consider the definition of slope here

OpenStudy (anonymous):

would it be -5 or -1/5

OpenStudy (uri):

Slope formula =(y2-y1)/(x2-x1)

OpenStudy (caitlinnr14):

slope is rise/run

OpenStudy (jdoe0001):

\(\large m=\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

So, in this case we want to subtract y1 from y2 (6-1) and x1 from x2 (4-5)

OpenStudy (anonymous):

so the answer is -5

OpenStudy (anonymous):

Then divide the change in Y by change in X

OpenStudy (anonymous):

Yes! You are correct

OpenStudy (anonymous):

thank you

OpenStudy (caitlinnr14):

answer is 1/5

OpenStudy (caitlinnr14):

right?

OpenStudy (anonymous):

Well, we have y2-y1, which is 6-1, or 5

OpenStudy (anonymous):

Therefore, change in Y= 5 (the rise, if you will)

OpenStudy (anonymous):

Now, x2-x1= 4-5, or -1 (the run)

OpenStudy (anonymous):

Rise over run then becomes 5/-1, or -5

OpenStudy (anonymous):

The answer is -5

OpenStudy (caitlinnr14):

oh negative 1 thx

OpenStudy (anonymous):

If you make a graph, this makes even more sense, as a slope of -5 is much more intense then a slope of -1/5, and the slope between the two points is rather steep

OpenStudy (caitlinnr14):

hey can you help me with this problem @VostheBoss321 The sum of three consecutive integers is 267. What is the largest integer? i got 88